karlieeeeeee:

Is (1, 9) a solution to this system of equations? y = 7x + 2 y = x + 8

Convert:

Bro-

snowflake0531:

do you know how to use the substitution method for doing this?

snowflake0531:

ohhhhh, you're YOU again lolll, the one that convert has already explained a gazillion times -_- oop

Convert:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @snowflake0531 ohhhhh, you're YOU again lolll, the one that convert has already explained a gazillion times -_- oop \(\color{#0cbb34}{\text{End of Quote}}\) LOL

snowflake0531:

eeeeeeee

snowflake0531:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @karlieeeeeee fu shut up bruh- \(\color{#0cbb34}{\text{End of Quote}}\) aw well, you shoulda learned it by now O-O, from Convert and the other people who has helped -_- but do you know how to use the substitution method?

karlieeeeeee:

exactly

snowflake0531:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @Convert \(\color{#0cbb34}{\text{Originally Posted by}}\) @snowflake0531 \(\color{#0cbb34}{\text{Originally Posted by}}\) @karlieeeeeee fu shut up bruh- \(\color{#0cbb34}{\text{End of Quote}}\) aw well, you shoulda learned it by now O-O, from Convert and the other people who has helped -_- but do you know how to use the substitution method? \(\color{#0cbb34}{\text{End of Quote}}\) Like she said she "doesn't have time" to do it and just wants the answer. 😂 \(\color{#0cbb34}{\text{End of Quote}}\) lolll i think knowing how to know it will be better, not just bc you should, but also well, so you don't have to post every single one -_-

snowflake0531:

but do you know how to use the substitution thingyyyyyyyyyyy

Convert:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @snowflake0531 \(\color{#0cbb34}{\text{Originally Posted by}}\) @Convert \(\color{#0cbb34}{\text{Originally Posted by}}\) @snowflake0531 \(\color{#0cbb34}{\text{Originally Posted by}}\) @karlieeeeeee fu shut up bruh- \(\color{#0cbb34}{\text{End of Quote}}\) aw well, you shoulda learned it by now O-O, from Convert and the other people who has helped -_- but do you know how to use the substitution method? \(\color{#0cbb34}{\text{End of Quote}}\) Like she said she "doesn't have time" to do it and just wants the answer. 😂 \(\color{#0cbb34}{\text{End of Quote}}\) lolll i think knowing how to know it will be better, not just bc you should, but also well, so you don't have to post every single one -_- \(\color{#0cbb34}{\text{End of Quote}}\) Yess exactly, and with final exams and stuff you need to know how to do stuff

carmelle:

We can use the substitution method here.. So, substitute `x` with `1` y = 7x + 2 y = x + 8 y = 7(1) + 2 y = (1) + 8 Now we solve, 7 + 2 = ? 1 + 8 = ? If the answers are the same then (1,9) is a valid solution. Hope this helps!!

karlieeeeeee:

thank you

carmelle:

\(\color{#0cbb34}{\text{Originally Posted by}}\) @karlieeeeeee thank you \(\color{#0cbb34}{\text{End of Quote}}\) mhm np